40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 31 20:06:22 2013, Evan, Update, Locking, PRM/PR2 cavity 
    Reply  Thu Jan 31 23:48:45 2013, Koji, Update, Locking, PRM/PR2 cavity 
       Reply  Fri Feb 1 09:33:11 2013, Jamie, Update, Locking, PRM/PR2 cavity 
          Reply  Fri Feb 1 15:12:53 2013, Jenne, Update, Locking, PRM/PR2 cavity 
             Reply  Fri Feb 1 19:55:33 2013, Jenne, Update, Locking, PRM/PR2 cavity 
                Reply  Fri Feb 1 23:12:42 2013, Koji, Update, Locking, PRM/PR2 cavity 
                Reply  Fri Feb 1 23:52:59 2013, rana, Update, Locking, PRM/PR2 cavity 
Message ID: 7985     Entry time: Fri Feb 1 15:12:53 2013     In reply to: 7981     Reply to this: 7986
Author: Jenne 
Type: Update 
Category: Locking 
Subject: PRM/PR2 cavity 

 After Jamie did all the work this morning on the POP table, I was able to get the cavity to lock.  It's not very stable until I engage the boost filters in the PRCL loop.  After locking, I tuned up the alignment a bit more.  Now we're taking mode scan data.  Look for results hopefully shortly after Journal Club!

ELOG V3.1.3-